OpenStudy (anonymous):

Help Please!! Write the equation 0.3x^ 2 + 5x - 7 = 0 in general form and then choose the value of "b." 70 50 5

OpenStudy (anonymous):

\[ax^2+bx+c=0\]\[.3x^2+5x-7=0\] what is \(b\)?

OpenStudy (anonymous):

i thought it was 5, but that was wrong,

OpenStudy (anonymous):

really????

OpenStudy (anonymous):

yeah,

OpenStudy (anonymous):

the only thing i can think they might want you to do is multiply by 10 and get \[3x^2+50x-70=0\]which would make \(b=50\)

OpenStudy (anonymous):

okay, thanks,

OpenStudy (anonymous):

that is my guess, i would have said 5 if you had not said it was wrong

OpenStudy (anonymous):

yeah, cause im doing online classes and it shows me which problems i get wrong

OpenStudy (anonymous):

http://media01.owotw.com/g_ima01_2009/7/17.gif